# Deeper quantum circuits via pseudo-twirling coherent errors mitigation in non-Clifford gates

### v1

@@ II.2 PST application alternatives
  • Circuit PST - Applying PST to the entire circuit, with the twirling gates positioned solely at the circuit’s edges.
@@
  • Intra-gate PST - Implementing PST within each gate.
  • Hybrid PST-RC - In Clifford gates, combine Intra-gate PST with RC, or RC with circuit PST
The first three options are illustrate in Fig. 2c. Given the absence of a clear cost analysis, it appears advantageous to consider all three levels of implementation, or at the very least, the first and third options. Implementing PST in shorter intervals (intra-gate PST) is an effective strategy to prevent the accumulation and growth of coherent error. Employing PST across the entire circuit is also beneficial, as it can mitigate the accumulation of errors resulting from non-local (non-gate-based) mechanisms.
@@
Let us elaborate on the intra-gate PST approach. Taking the example of the cross-resonance CNOT gate [25] depicted in Fig. 4a, each non-Clifford element (pulse) can be pseudo-twirled independently within the PST framework as shown in Fig. 4b. This contrasts with RC, which cannot be applied to the non-Clifford elements of the gate. Additionally, each one of two segments of the cross-resonance pulse can be divided into multiple shorter pulses, allowing for the application of PST across these smaller segments (Fig. 4c). The prospective advantage here is to mitigate coherent errors before they accumulate. More detailed results are available in the Preliminary Results section.
In the context of Clifford gates such as the CNOT, conventional RC can be applied at the gate level, while employing PST within the gate’s internal elements. RC aids in addressing calibration errors that PST does not address. Meanwhile, PST’s role is to decrease the level of coherent errors experienced by RC, resulting in reduced incoherent errors and less fluctuations in the post-processing averaging.
@@
Figure 4: (a) A schematic representation of the pulse schedule that generates the echo cross-resonance gate as used by IBM. (b) Intra-gate PST: since PST is not limited to Clifford gates, it can be applied separately to each of the XZ pulses. (c) Taking it a step further, the possibility arises to decompose the π/4 XZ pulse into multiple shorter pulses and apply PST to each of these pulses. In the preliminary results section, it is demonstrated that this splitting enhances the effectiveness of PST in reducing coherent errors.
II.3 Integration and Compatibility with QEM
@@
It is evident that PST, by itself, will only provide a limited solution to errors in quantum devices, as it converts coherent errors into incoherent errors, which still need to be addressed using QEM. Therefore, the impact of PST largely depends on its compatibility with QEM methods. While some error mitigation protocols are expected to be compatible with PST, others that rely on the twirled noise to be a Pauli channel, such as PEC and PEA, may not be. In what follows we demonstrate numerically and experimentally that PST is not only compatible with the Adaptive KIK QEM method [22] it actually improve the resilience of the method to crosstalk.
III Numerical Results
III.1 Intra-Gate PST
@@
As alluded to in Fig. 4c, implementing PST on smaller segments of a pulse allows for further improvement in coherent error suppression. To illustrate this point, we conducted a simulation involving the slicing of a $\pi/2$ cross-resonance pulse. We applied PST to each slice and plotted the error in the entire evolution compared to the ideal $\pi/2$ cross-resonance pulse. The results, shown in Fig. 5, reveal that error suppression enhances as the number of slices increases. We observed that the level of suppression is inversely proportional to the number of slices (indicated by the orange curve). Intuitively, since PST suppress the leading order of the coherent error it becomes more effective when applied to thiner slices where the higher order terms (of the coherent error) are negligible. Thus, in thin slices the PST mitigates the coherent error before it has chance to accumulate.
@@
Clearly, there comes a point where the length of the slice becomes comparable to that of the PST Pauli operators, and further slicing will not be possible. Additionally, residual coherent errors in the PST Pauli operators may also set a performance limit. Nonetheless, our results suggest that before reaching these limiting factor intra-gate PST improves the coherent error suppression.
Refer to caption Figure 5: Numerical simulation illustrating the impact of intra-gate PST on coherent errors. In this example, the ideal evolution corresponds to the cross-resonance pulse utilized in IBM’s implementation of the CNOT gate. Slicing the pulses into multiple shorter segments and applying PST to each of them results in a reduction of the final coherent error. The x-axis represents the number of slices (pulses), while the y-axis indicates the trace norm distance from the ideal pulse. The solid orange line corresponds to a fit using the equation $\alpha$/(number of slices), where $\alpha$ is the fitting parameter.
III.2 PST as a “Hermitianizer” of the Error Channel
To illustrate the effect of PST on incoherent errors, we consider the same cross-resonance pulse slicing example shown in Fig. 5, but this time with coherent errors replaced by incoherent errors. Figure 6 displays the non-hermiticity measure $\frac{1}{2}\left\|N-N^{\dagger}\right\|/\left\|N\right\|$ of the error channel as a function of the number of slices. This measure equals one for anti-Hermitian operators and zero for Hermitian operators. The sharp drop of the curve illustrates our analytical finding that PST makes the error channel more Hermitian. We have confirmed this in different circuits and with various error models, including coherent errors.
@@ III.2 PST as a “Hermitianizer” of the Error Channel
Figure 6: The effect of PST on incoherent errors (noise). The setup is the same cross-resonance pulse setup presented in Fig. 5, but this time with incoherent error instead coherent error. The $y$ axis shows the non-hermiticity measure $\frac{1}{2}\left\|N-N^{\dagger}\right\|/\left\|N\right\|$ of the noise channel $N$ as function of the number of slices. The fast drop illustrates our analytical finding that PST makes the noise channel more hermitian.
@@
Figure 7: Integration of pseudo-twirling with the Adaptive KIK quantum error mitigation method. Different x values represent the trace norm error when implementing a 3-qubit transverse Ising time evolution. The error increase in the x axis in obtained by increasing the coherent error amplitude. The y-axis demonstrates errors after the application of various protocols. The dashed line represents errors without any techniques applied. The green curve corresponds to the first-order Adaptive KIK mitigation without PST and the red curve depicts the error reduction when PST is applied without KIK. The combination of KIK and PST (blue curve) showcases a significant error suppression. The inset displays the same plot in a logarithmic scale. Our goal is to explore this powerful integration of PST and quantum error mitigation, both analytically and through cloud-based experimental platforms.
III.3 Integration of PST and Quantum Error Mitigation
@@
As explained earlier, PST alone will not suffice. Ideally, it should completely eliminate coherent errors and transform the noise into a fully Hermitian noise channel. However, the resulting noise channel can still be too strong to extract meaningful results from a given quantum computer. Therefore, PST must be followed by a quantum error mitigation (QEM) protocol. We chose the ‘Adaptive KIK’ QEM method [22] due to its is operational simplicity and its compatibility with non-Clifford gates. To exemplify the compatibility of PST with this QEM method, we conducted a three-qubit transverse Ising model simulation with both coherent and incoherent errors. Figure 7 illustrates that PST significantly enhances Adaptive-KIK performance, as evident from the shift from the green curve to the blue curve. To emphasize the insufficiency of PST alone, we have also plotted in red the error when PST is applied without KIK.
IV Experimental Results
@@
To illustrate the utility of PST to gate-calibration we run a ‘deep calibration’ experiment where we calibrate a sequence of 81 $\pi/2$ ZX echo cross resonance pulses. Such a large number of repetitions amplifies the smallest coherent errors. Yet, a trivial execution of such a long circuit is prone to substantial incoherent errors (noise) that will dramatically effect the calibration curve and lead to an erroneous calibration as demonstrated in [22]. Thus, we use the Adaptive KIK [22] QEM methodology for removing the incoherent errors. We measure the probability to measure the system in the initial $\left|00\right\rangle$ (survival probability) after 81 sequences for different stretch factor $\chi$ of the CR pulse amplitude. A value of one corresponds to the default amplitude value provided by IBM.
@@
By comparing the results of applying KIK with and without PST we observe in Fig. 8 that i) with PST (Fig. 8b) the KIK converges faster (less mitigation orders are needed) compared to the same experiment without PST (Fig. 8a). ii) The value of the calibrated amplitude $\chi(y=1/2)$ is affected by the PST. iii) Without PST for some amplitudes we get unphysical values. to the best of our knowledge, such non-physical values can occur for two reasons: 1) the mitigation order is too small with respect to the window size parameter in the KIK method (see simulations in the appendix of [22]) and the other reason is non-Markovian effects. One of the sources of non-Markovian effect in the present setup is crosstalk interaction with adjacent qubits. If the order is too low we expect the unphysical values to disappear as the order increases. Indeed $n=3$ appear to be more physical than $n=2$, yet, since the two result differ convergence is not observed and we cannot confirm that unphysical values disappear in higher orders. Moreover, when the mitigation order is too low PST is not expected to help. The fact that PST does removes the unphysical values and speed up convergence suggest that that the unphysical value without PST are related to non-Markovian effects. We point out that in other experiments we observed this behavior when neighboring idle qubits are not protected using dynamical decoupling.
@@
Note that what we control is $\chi$ the actual drive of the ZX term $\chi\alpha\cos(\phi)$. In PST calibration it is not needed to know $\phi$ explicitly, $\chi$ is scanned until the survival probability obtain the value of $1/2$ which means that $\chi\alpha\cos(\phi)=\pi/4$ (the value that corresponds to $\pi/2$ CR pulse).
@@
Figure 8: A ‘deep calibration’ experiment of a cross resonance gate is conducted. The initial state is set to $\left|00\right\rangle$, then a sequence of 81 $\pi/2$ echo cross resonance pulses are executed, aiming at amplifying small coherent errors. The x-axis describes the cross resonance amplitude stretch factor $\chi$.. At such depths, the incoherent errors become pronounced, and error mitigation has to be applied. n=0 corresponds to no error mitigation, and n>0 corresponds to higher-order Adaptive KIK error mitigation methods. The point $\chi(y=0.5)$, where the graph intersects the value $y=0.5$ (vertical lines), sets the stretch factor that should be used to recalibrate the amplitude of the cross resonance pulse. The left figure (a) shows the results without PST, and the right figure (b) shows the results with PST. Not only does PST speed up the convergence (n=2 suffices with PST), but PST also prevents the unphysical result observed without PST (red regime, probability larger than one). These non-physical results can arise due to non-Markovian effects, which PST suppresses. Although the vertical lines indicate that the amplitude should be modified only by $0.2$ with respect to the default value set by IBM ($\chi=1$), it is observed that the error in the survival probability is as big as $0.1$ if $\chi=1$ is used instead of $\chi(y=1/2)\simeq 1.0021$.
@@
One might worry that for $\phi\neq 0$ a residual ZY term will remain and generate a coherent error. However as explained in is Sec., and demonstrated experimentally next, PST eliminated the small $\sin\phi\sigma_{z}\otimes\sigma_{y}$ terms. To experimentally demonstrate this point we carry out a variant of the previous experiment where a $\pi/2$ single-qubit $y$ rotation is added to the controlled qubit (in $ZX$ interaction the right qubit is the controlled qubit) just before the measurement. This makes the survival probability sensitive to $ZY$ rather than to $ZX$. Without PST it is expected that the outcomes will show some dependence on the amplitude since $\sin\phi\sigma_{z}\otimes\sigma_{y}$ is not exactly zero if $\phi$ is not perfectly calibrated to zero. As show in Fig. 9b after applying PST and KIK we find that calibration curve is flat as expected. As a comparison we show the result without PST on 9a. The mean square error $\sqrt{SSD}$ in fitting a straight line to the data is presented in the yellow boxes.
@@
Figure 9: The same experiment as shown in Fig. 8, but this time, instead of measuring the population of the state $\left|00\right\rangle\left\langle 00\right|$, we measure the probability of populating the state $\left|0,-x\right\rangle\left\langle 0,-x\right|$. This is achieved by applying a $R_{y}(\pi/2)$ rotation before measuring the controlled qubit. Consequently, the outcome is sensitive to the ZY term instead of the ZX term. Since PST suppresses the residual ZY term and other coherent errors as well, we expect the calibration curve to be flat as the amplitude stretch is varied. As shown in Fig. 9b, the experimental results validate this prediction. In contrast, the n=3 curve without PST is not flat. To see this more quantitatively, we calculate the $\sqrt{SSD}$ score (SSD - Sum of Squared Differences) that indicates the error in a fit to a straight line, and find that it is ten times smaller with PST compared to without PST.
@@
Finally we point out an important observation regarding the use of virtual Z gates in the pseudo twirling gates. Due to their simplicity and very low error rates, virtual Z gates are very popular in quantum computers setups. In short, the Z rotation can pushed further dow the circuit by modifying the phases of pulses that generate gates that do not commute with Z. However, despite their advantage, as explained in Appendix II , virtual gate are not suitable for PST. Thus, in our experiment we implement the $Y$ gate by changing the phase of the $X$ pulse and not by implementing $XZ_{v}$ (where where $Z_{v}$ is a virtual Z gate. For $Z$ gate we use $Z_{p}=Y_{p}X$ or $XY_{p}$. When the noise is very weak these two options are equivalent. We show that using virtual gates make twirling that are suppose to be different become the same.
V Summary and concluding remarks

### v2

@@ III.2 Alternative Ways for Applying PST in a Circuit
  • Edges PST - Applying PST to the entire circuit, with the twirling gates positioned solely at the circuit’s edges.
@@
  • Intra-gate PST - Implementing PST on elements within each gate.
  • Hybrid PST-RC - In Clifford gates, combine Intra-gate PST with RC, or RC with edges PST.
The first three options are illustrated in Fig. 2c. In principle, it appears advantageous to consider all three levels of implementation, or at the very least, the first and third options. Implementing PST in shorter intervals (intra-gate PST) is an effective strategy to prevent the accumulation of coherent errors. Employing PST across the entire circuit is also beneficial, as it can mitigate the accumulation of errors resulting from non-local (non-gate-based) mechanisms.
@@
Let us elaborate on the intra-gate PST approach. Taking the example of the cross-resonance CNOT gate [29] depicted in Fig. 3a, each non-Clifford element (pulse) can be pseudo twirled independently within the PST framework as shown in Fig. 3b. This contrasts with RC, which cannot be applied to the non-Clifford elements of the gate. Additionally, each one of two segments of the cross-resonance pulse can be divided into multiple shorter pulses, allowing for the application of PST across these smaller segments (Fig. 3c). The prospective advantage here is to mitigate coherent errors before they accumulate. To quantify this claim we show in Appendix III that when the number of slices $m$ is sufficiently large so that neither the drive nor the coherent error changes on the scale of a single slice, the evolution operator is given by:
$\mathcal{K}_{PST}^{(m)}=\mathcal{U}_{T}e^{\frac{T}{m}\intop\mathcal{U}^{\dagger}(t)\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{L}(P_{\alpha}H_{coh}P_{\alpha})\mathcal{U}(t)dt}.$ (22)
While the integral is $m$ independent, the coefficient before it scales like $1/m$ which agrees with the numerical observations in Sec. IV.1. We can quantify the residual noise using the operator norm (see Appendix IV for the explanation why this norm is used)
$\left\|\mathcal{K}_{PST}^{(m)}-\mathcal{U}_{T}\right\|_{op}\simeq\frac{T}{m}\left\|\intop\mathcal{U}^{\dagger}(t)\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{L}(\mathcal{P_{\alpha}}\mathcal{H}_{coh}\mathcal{P_{\alpha}})\mathcal{U}(t)dt\right\|_{op},$ (23)
which clearly scales like $1/m$.
In the context of Clifford gates such as the CNOT, conventional RC can be applied at the gate level, while employing PST within the gate’s internal elements. As PST mitigates the non-rotational coherent errors, RC will have to cope mostly with the residual rotation errors.
@@
Figure 3: (a) A schematic representation of the pulse schedule that generates the IBM echo cross-resonance gate. (b) Intra-gate PST: since PST is not limited to Clifford gates, it can be applied separately to each of the ZX pulses (red lines). (c) Taking it a step further, it is possible to decompose each of the π/4 ZX pulses into $m$ shorter pulses and apply PST to each of these shorter pulses. (d) The slicing effect on a $\pi/2$ ZX pulse. The x-axis represents the number of slices (pulses), while the y-axis indicates the operator norm error from the ideal pulse $\left\|\mathcal{K}_{PST}^{(m)}-\mathcal{U}_{T}\right\|_{op}$. The solid orange line corresponds to a fit using the function $b/m$, where $b$ is a fitting parameter.
III.3 Integration and Compatibility with Quantum Error Mitigation
@@
It is evident that PST, by itself, will only provide a limited solution to errors in quantum devices, as it converts coherent errors into incoherent errors, which still need to be addressed using QEM. Therefore, the impact of PST largely depends on its compatibility with QEM methods. PST was mainly introduced for handling non-Clifford gates. Yet QEM methods that are restricted to Cliffords gates such PEC and PEA may still benefit from intra-gate PST and circuit level PST.
Finally, we point out that much like the ‘Adaptive KIK’ QEM method [24], the feasibility of executing PST hinges on the ability to control the signs of the different terms in the driving Hamiltonian. We have shown in Section II.3 and II.4 that PST turns the non-rotation coherent errors and non Hermitian dissipators into a Hermitian Lindblad dissipator. Hence it is expected that PST will be compatible with the Adaptive KIK method. In Sec. IV.3 and V, we demonstrate numerically and experimentally that PST is not only compatible with Adaptive KIK, but substantially improves the resilience of the method to crosstalk.
IV Numerical Results
IV.1 Intra-Gate PST
@@
As alluded to in Fig. 3c, implementing PST on smaller segments of a pulse allows for further improvement in coherent error suppression. To illustrate this point, we conducted a simulation involving the slicing of a $\pi/2$ cross-resonance pulse. The uncontrolled coherent error is a $\sigma_{z}\otimes\sigma_{z}$ term that accompanies the $\sigma_{z}\otimes\sigma_{x}$ driving Hamiltonian. We have applied PST to each one of the $m$ slices and plotted the error in the resulting evolution operator $\mathcal{K}_{PST}^{(m)}$ with respect to the ideal evolution $\mathcal{U}_{ideal}$ of a $\pi/2$ cross-resonance pulse. Figure 3d shows that the resulting error, which we quantify using $\left\|\mathcal{K}_{PST}^{(m)}-\mathcal{U}_{ideal}\right\|_{op}$, reduces as the number of slices increases. In agreement with (22) we observe that the level of suppression is inversely proportional to the number of slices. The orange curve is a fit of the function $b/m$ to the data, where $b$ is a fit parameter that corresponds to the norm operator of the first Magnus term of the noise [see (23)]. Intuitively, in thin slices the PST mitigates the coherent error before it has a chance to accumulate.
@@
In practice, at some point the duration of the slice becomes comparable to that of the PST Pauli operators, and the small error in these operators sets a performance limit. Nonetheless, our results suggest that before reaching this limiting factor intra-gate PST improves the coherent error suppression.
IV.2 PST as a “Hermitianizer” of the Error Channel
To illustrate the effect of PST on incoherent errors, we consider again a $\pi/2$ echo cross-resonance pulse, but this time with coherent errors replaced by incoherent errors. A $T_{1}$ noise (amplitude damping) is applied to each qubit. To quantify the non-Hermiticity of a noise channel $\mathcal{N}=\mathcal{U}^{\dagger}\mathcal{K}$ we use the measure $G=\left\|\mathcal{N}-\mathcal{N}^{\dagger}\right\|_{op}/\left\|2\mathcal{N}\right\|_{op}$. $G$ equals one for anti-Hermitian operators, and zero for Hermitian operators. Figure 4 displays how the PST suppresses $G$ for different level of initial non-Hermiticity. The initial non-Hermiticity was controlled by varying the strength of the amplitude damping noise. As a result, the Hermitian part of the noise increases as well. We have observed the same behavior when using randomly chosen Lindbladians from an ensemble of random Lindbladians [36].
@@ IV.2 PST as a “Hermitianizer” of the Error Channel
Figure 4: The effect of PST on incoherent errors (noise). The setup is the same as the cross-resonance pulse setup presented in Fig. 3, but this time with incoherent errors instead coherent errors. The $y$ axis shows the non-Hermiticity suppression factor $G_{no\>PST}/G_{PST}$ where $G=\frac{1}{2}\left\|\mathcal{N}-\mathcal{N}^{\dagger}\right\|_{op}/\left\|\mathcal{N}\right\|_{op}$ measure the non-Hermiticity of the noise channel $\mathcal{N}$.
@@
Figure 5: (a) Integration of pseudo twirling with the Adaptive KIK quantum error mitigation method. Different $x$ values represent the operator norm error of the evolution operator when implementing a 3-qubit transverse Ising time-evolution. The error increase in the $x$ axis is obtained by increasing the coherent error amplitude. The $y$-axis shows the operator norm error after the application of various protocols. The dashed line represents errors without any technique applied. This is a ‘break-even’ line where any result below it represents an improvement. The green curve corresponds to the first-order Adaptive KIK mitigation without PST and the red curve depicts the error reduction when PST is applied without KIK. The combination of KIK and PST (blue curve) showcases a significant error suppression. The inset displays the same plot in a logarithmic scale. For each data point we have used twenty twirlings. (b) Using a hundred twirlings, for initial coherent error of 0.25 (corresponding to the right edge of Fig. (a)) substantially improves the error suppression of KIK+PST as shown by the blue histogram.
IV.3 PST+QEM Integration in the Ising Model
@@
As mentioned earlier, PST alone provides only a partial solution to errors in quantum computers. Ideally, it transforms the noise and coherent errors into a Hermitian noise channel. However, the resulting noise channel can still be too strong to extract meaningful results from a given quantum computer. Therefore, PST must be complemented by a quantum error mitigation protocol. We choose the ‘Adaptive KIK’ QEM method [24] due to its operational simplicity and its compatibility with non-Clifford gates. To exemplify the compatibility of PST with this QEM method, we conducted a three-qubit transverse Ising model simulation with both coherent and incoherent errors. The coherent error Hamiltonian was chosen to be $\{\sigma_{z,i}\}$. Such an error can arise from imprecise determination of the qubits frequencies. Finally, the evolution before the PST is given by:
$\mathcal{K}=e^{-i\,J\sum_{i=1}^{2}\mathcal{H}_{z_{i},z_{i+1}}-i\,g\sum_{i=1}^{3}\mathcal{H}_{x_{i}}-i\,\epsilon\sum_{i=1}^{3}\mathcal{H}_{z_{i}}+\mathcal{L}},$ (24)
where $J$ and $g$ are parameters of the transverse Ising Hamiltonian and $\epsilon$ sets the amplitude of the coherent error. The pulse inverse $\mathcal{K}_{I}$ needed for the KIK protocols is
$\mathcal{K}_{I}=e^{+i\,J\sum_{i=1}^{2}\mathcal{H}_{z_{i},z_{i+1}}+i\,g\sum_{i=1}^{3}\mathcal{H}_{x_{i}}-i\,\epsilon\sum_{i=1}^{3}\mathcal{H}_{z_{i}}+\mathcal{L}}.$ (25)
Importantly, in the pulse inverse $\mathcal{K}_{I}$ the sign of $\epsilon$ remains the same since it represents an uncontrollable coherent error that does not depend on the external drive. The first order KIK mitigated operator is
$\mathcal{K}^{(1)}=\frac{3}{2}\mathcal{K}^{PST}-\frac{1}{2}\mathcal{K}^{PST}\mathcal{K}_{I}^{PST}\mathcal{K}^{PST}.$ (26)
Figure 5(a) illustrates that PST significantly enhances the KIK performance, as evident from the shift from the green curve to the blue curve. To emphasize the insufficiency of PST alone, we have also plotted in red the error when PST is applied without KIK. At the rightmost point the PST+KIK reduces the error by a factor of 65.
In this simulation we randomly sampled the twirling space and used only 20 twirling instances. In the implementation of the evolution $\mathcal{K}$ and $\mathcal{K}\mathcal{K}_{I}\mathcal{K}$ needed for the KIK protocol we used the same twirlings for all $\mathcal{K}$ and $\mathcal{K}_{I}$. Figure 5(b) shows the dependence of the error suppression ratio of the PST combined with 1st order KIK protocol with respect to the original error for $n_{PST}=20$ twirlings (orange) and $n_{PST}=100$ twirlings (blue) . The original noise level is 0.25 (right hand side of Fig. 5(a)). These histograms were generated by repeating the simulation 800 times with different random twirlings. For the simulation parameter we used $J=0.1,g=0.2,\xi=0.00465$. $\epsilon$ was varied from $0.005$ to $0.05$. For the dissipative term $\mathcal{L}$ we use $T_{1}$ noise (amplitude damping) for the different qubits such that $\mathcal{L}=0.5\mathcal{L}_{1}^{T_{1}}+1.7\mathcal{L}_{2}^{T_{1}}+0.3\mathcal{L}_{3}^{T_{1}}$ where the subscript designates the qubit index. The initial state is $\left|000\right\rangle$. The the error was quantified using the operator norm as in Fig. 5.
V Experimental Results
@@
To illustrate the utility of PST to gate-calibration we run a ‘deep calibration’ experiment where we calibrate a sequence of 81 echo cross-resonance (ZX) $\pi/2$ pulses. Such a large number of repetitions amplifies the smallest coherent errors. Yet, a trivial execution of such a long circuit is prone to substantial incoherent errors (noise) that will dramatically effect the calibration curve and lead to an erroneous calibration as demonstrated in [24]. Thus, we use the Adaptive KIK [24] QEM methodology for removing the incoherent errors. We measure the probability to find the system in the initial $\left|00\right\rangle$ state (survival probability, SP) after the 81 repetitions sequence for different stretch factor $\chi$ of the cross-resonance pulse amplitude. $\chi=1$ corresponds to the default amplitude value provided by IBM. In this experiment we use the ibmq_kolkata quantum computer.
@@
By comparing the results of applying KIK with and without PST we observe in Fig. 6 that i) with PST (Fig. 6b) the KIK converges faster (less mitigation orders are needed) compared to the same experiment without PST (Fig. 6a). ii) The value of the calibrated amplitude $\chi(SP=1/2)$ is affected by the PST. iii) Without PST for some $\chi$ values we observe unphysical values. To the best of our knowledge, such non-physical values can occur for two reasons. First, if the mitigation order is too small with respect to the window size parameter in the KIK method (see simulations in the appendix of [24]) nonphysical values can appear. Yet, in this case, it is expected that the unphysical values will disappear in higher mitigation orders. The other reason is pronounced non-Markovian effects. One of the sources of non-Markovian effects in the present setup is crosstalk interaction with adjacent qubits. $n=3$ appears to be more physical than $n=2$, yet, since the two results differ, convergence is not observed and we cannot confirm that unphysical values disappear in higher orders. Moreover, when the mitigation order is too low PST is not expected to help. The fact that PST removes the unphysical values and speeds up convergence suggests that the unphysical values without PST are related to non-Markovian effects. We point out that in other circuits we have suppressed unphysical values without PST by applying dynamical decoupling to neighboring idle qubits.
@@
While we control $\chi$, the actual drive of the ZX term is $\chi a\cos(\phi)$ where $\phi$ is the phase of the drive. In PST calibration it is not needed to know $\phi$ explicitly, $\chi$ is scanned until the survival probability obtain the value of $1/2$ which means that $\chi a\cos(\phi)=\pi/4$ (the value that corresponds to $\pi/2$ cross-resonance pulse).
@@
Figure 6: A ‘deep calibration’ experiment of a cross-resonance gate is conducted on ibmq_kolkata. The initial state is set to $\left|00\right\rangle$, then a sequence of 81 $\pi/2$ echo cross-resonance pulses is executed, aiming at amplifying small coherent errors. The $x$ axis describes the cross-resonance amplitude stretch factor $\chi$. At such depths, the incoherent errors become pronounced, and error mitigation has to be applied. $n=0$ corresponds to no error mitigation, and $n>0$ corresponds to higher-order Adaptive KIK error mitigation methods. The point $\chi(SP=0.5)$, where the graph intersects the value $SP=0.5$ (vertical lines), sets the stretch factor that should be used to recalibrate the amplitude of the cross-resonance pulse. The left figure (a) shows the results without PST, and the right figure (b) shows the results with PST using twenty random twirlings. Not only does PST speeds up the convergence ($n=2$ suffices with PST), but PST also prevents the unphysical results observed without PST (shaded-purple region, probability larger than one). These non-physical results can arise due to non-Markovian crosstalk effects, which PST suppresses. Although the vertical lines indicate that the amplitude should be modified only by $0.2\%$ with respect to the default value set by IBM ($\chi=1$), it is observed that the error in the survival probability is as big as $0.1$ if $\chi=1$ is used instead of $\chi(SP=1/2)\simeq 1.0021$.
@@
One might worry that for $\phi\neq 0$ a residual ZY term will remain and generate a coherent error. However as explained in Sec. III.1, PST eliminates the small $a\sin\phi\,\sigma_{z}\otimes\sigma_{y}$ terms. To experimentally demonstrate this point we carry out a variant of the previous experiment where a $\pi/2$ single-qubit $y$ rotation is added to the target qubit just before the measurement. This makes the survival probability sensitive to $ZY$ rather than to $ZX$. Without PST, it is expected that the outcomes will show some dependence on the amplitude since $\chi a\sin\phi\,\sigma_{z}\otimes\sigma_{y}$ is not exactly zero if $\phi\neq 0$. This behavior is observed in 7a for different mitigation order $n$. However, Fig. 7b demonstrates that after applying PST and KIK we find that the calibration curve is flat as expected. The root of the sum square residuals $\sqrt{SSR}$ in fitting a straight line to the data is presented in the yellow boxes. For $n=3$ the $\sqrt{SSR}$ score with PST is an order of magnitude smaller compared to the no PST case. In this last experiment we used the Taylor coefficients in the KIK method in order to reduce the error bars. Furthermore, for better visibility, in this experiment the sequence contained 41 repetitions of $\pi/2$ echo cross-resonance pulses.
The error of the means $\sigma$ used for plotting the error bars in the PST plot in Fig. 8 and Fig. 9 were calculated using
$\sigma^{2}=\frac{1}{n_{PST}-1}\sum_{j=1}^{n_{PST}}(R_{j}-\frac{1}{n_{PST}}\sum_{k=1}^{n_{PST}}R_{k})^{2},$ (27)
where $n_{PST}$ is the number of PST realizations used, and $R_{j}$ is the mitigated survival probability calculated by averaging over all the shots in a given realizations. At first, it may seem counter intuitive that this simple formula captures correctly the two components of the variance: i) the variance in the measurement of each $R_{k}$ that is determined by the number of shots in each realization, and ii) the variance in the values of $R_{j}$ with respect to its mean ($\frac{1}{n_{PST}}\sum_{k=1}^{n_{PST}}R_{k}$) that is related to the PST twirling of the coherent error. Such a separation to the variance of means and means of variances is possible provided that a good estimation of the means is known [37]. However, our measurements do not provide the exact means but the ‘noisy means’ that includes the additional uncertainty generated by the finite number of shots. We numerically corroborate that the variance of noisy means formula (27) agrees with the variance observed in simulation.
@@
Figure 7: The same experiment as shown in Fig. 6, but this time, instead of measuring the population of the state $\left|00\right\rangle$, we measure the probability of populating the state $\left|0,-x\right\rangle$. This is achieved by applying a $R_{y}(\pi/2)$ rotation before measuring the target qubit. Consequently, the outcome is sensitive to the ZY term instead of the ZX term. Since PST suppresses the residual ZY term and other coherent errors as well, we expect the calibration curve to be flat as the amplitude stretch is varied. As shown in Fig. 7b, the experimental results validate this prediction. In contrast, the $n=3$ curve without PST is not flat. To see this more quantitatively, we calculate the $\sqrt{SSR}$ score (SSR - Sum of Squared Residuals) that indicates the error in a fit to a straight line, and find that it is ten times smaller with PST compared to the same experiment without PST. For clarity, the fitted straight lines are not shown. We used forty random twirling in this experiment.
@@
Finally we point out an important observation regarding the use of virtual Z gates in the pseudo twirling Paulis. Due to their simplicity and zero error rates, virtual Z gates [38] are very popular in quantum computers setups. In short, Z rotations can be pushed further down the circuit by modifying the phases of pulses that generate gates that do not commute with Z. However, despite their advantage, as explained in Appendix V, a virtual Z gate is not suitable for generating the Z gates needed for PST. Essentially, it makes twirlings that should be different become the same, effectively reducing the space of available twirlings. Thus, in our experiment, we implement the $Y_{pulse}$ gate by changing the phase of the $X$ pulse and not by implementing $XZ_{v}$ (where $Z_{v}$ is a virtual Z gate). For $Z$ gates we use $Z_{pulse}=Y_{pulse}X$ or $XY_{pulse}$. See also a similar effect of virtual Z on dynamical decoupling experiments [39].
VI Summary and concluding remarks